Fill in the blanks to make the following statements true:
(– 4) + [15 + (–3)] = [– 4 + 15] +............

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let us assume the missing integer be x,
Then,
= (–4) + [15 + (–3)] = [–4 + 15] + x
= (–4) + [15–3)] = [–4 + 15] + x
= (-4) + [12] = [11] + x
= 8 = 11 + x
By sending 11 from RHS to LHS it becomes -11,
= 8–11 = x
= x = -3
Now substitute the x value in the blank place,
= (–4) + [15 + (–3)]
= [–4 + 15] + -3 ... [This equation is in the form of Associativeproperty of Addition]

Thanks